A candle burns at a rate of 1 inch per hour. The original height of the candle is 12 inches. The function of the candle burning can be represented by f(x)=-x+12. Find x when f(x)=8.
Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

f(x)=-x+12. Find x when f(x)=8
-------------------
Wherever there's x, sub 8.

Answer by stanbon(75887)

A candle burns at a rate of 1 inch per hour. The original height of the candle is 12 inches. The function of the candle burning can be represented by f(x)=-x+12. Find x when f(x)=8.
---
Solve:: -x+12 = 8
-x = -4
x = 4
----------------
Cheers,
Stan H.
-------------